Learned Corporation has provided the following information:

Cost per Unit Cost per Period
Direct materials $ 5.60
Direct labor $ 3.90
Variable manufacturing overhead $ 1.50
Fixed manufacturing overhead $ 22,500
Sales commissions $ 0.50
Variable administrative expense $ 0.40
Fixed selling and administrative expense $ 7,500

Required:

a. For financial reporting purposes, what is the total amount of product costs incurred to make 5,000 units?
b. For financial reporting purposes, what is the total amount of period costs incurred to sell 5,000 units?
c. If the selling price is $23.10 per unit, what is the contribution margin per unit sold? (Round your answer to 2 decimal places.)
d. If 6,000 units are produced, what is the total amount of direct manufacturing cost incurred?
e. If 6,000 units are produced, what is the total amount of indirect manufacturing costs incurred?
Answer #1

a) Total amount of product cost

Direct material (5000*5.6) 28000
Direct labor 19500
Variable manufacturing overhead (5000*1.5 7500
Fixed manufacturing overhead 22500
Total product cost 77500

b) Total period cost = (5000*.90+7500) = 12000

c) Contribution margin per unit = 23.10-11.90 = 11.20

d) Total amount of direct manufacturing cost = (5.6+3.9)*6000 = 57000

e) Total amount of indirect manufacturing cost = (6000*1.5+22500) = 31500